Suppose you just purchased a bond with 15 years to maturity that pays an annual coupon of $20.00 and is selling at par. Calculate the one-year holding period return for each of these two cases:

A) The yield to maturity is 3.50% one year from now

B) The yield to maturity is 1.50% one year from now
Answer #1

a.

Calculating Bond Price after 1 year,

Using TVM calculation,

PV = [FV = 1,000, PMT = 20, N = 14, I = 0.035]

PV = $836.19

Return Rate = (836.19 - 1,000)/1,000 = -16.38%

b.

Calculating Bond Price after 1 year,

Using TVM calculation,

PV = [FV = 1,000, PMT = 20, N = 14, I = 0.015]

PV = $1,062.72

Return Rate = (1,062.72 - 1,000)/1,000 = 6.27%